OpenStudy (anonymous):

Condense The Expression \[2\ln x-3 lny\] Condense The Expression \[2\ln x-3 lny\] @Mathematics

OpenStudy (anonymous):

Try using the concepts that a*ln(b)=ln(b^a) and that ln(a)-ln(b)=ln(a/b)

OpenStudy (anonymous):

ln(x^2/y^3)